OpenStudy (anonymous):

PLEASE help me!! I am desperate! I do award a medal to whoever helps me. I just need someone to explain this to me. 2.Given the function k(x) = x2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: •k(x + c) •k(x) + c •k(cx) •c • k(x)

OpenStudy (anonymous):

Please help me understand this stuff.

OpenStudy (jdoe0001):

y = A (Bx + C ) + D C = vertical shrink, by a factor of "C" B = vertical shrink, by a factor of "B" D = vertical shift, D > 0, up, D < 0, down

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=3Q5Sy034fok

OpenStudy (jdoe0001):

if the original function was k(x) = x^2 then \(\bf k(x) = x^2\\ \quad \\ \quad \\ k(x) + c\implies x^2+c \qquad \textit{vertical shift upwards by "c"}\\ k(cx)\implies (c\cdot x)^2\implies c^2\cdot x^2 \qquad \textit{vertical shrink by a factor of }c^2\\ c\cdot k(x) \implies c\cdot x^2 \qquad \textit{vertical shrink by a factor of "c"}\)

OpenStudy (anonymous):

Thank you so much @jdoe0001! That helped me a lot!

OpenStudy (jdoe0001):

yw